If one ball is drawn without replacement from an urn with two red balls, three white balls, and three blue balls, what is the probability of drawing a red ball?
Thanks,
Mo

Answer by checkley75(3666)

SUCCESS/TOTAL OPTIONS=PROBABILITY
2/(2+3+3)=2/8=.25 OR 25% OF SUCCESS TO DRAW A RED BALL.